OpenStudy (anonymous):

How do you find the slope of 3x - 6y = -18

OpenStudy (hockeychick23):

get y by itself on one side: -6y= -3x-18 now divide both sides by -6 y= 1/2x+3 slope= 1/2

OpenStudy (studygurl14):

First step: Rearrange the equation into slope-intercept form so that you can easily pick out the slope. slope-intercept form is y=mx + b, where m = slope

OpenStudy (anonymous):